find the area of a triangle with vertices E(1,3) F(-4,-3) G(5,-4)


1
Expert's answer
2021-02-02T06:20:55-0500

Require to find the area of the triangle: "E(1,3),F(-4,-3),G(5,-4)"

Let "(x_{1},y_{1})=(1,3),(x_{2},y_{2})=(-4,-3),(x_{3},y_{3})=(5,-4)"

Area of the triangle "EFG" "=\frac{1}{2}\left [ x_{1}(y_{2}-y_{3})+x_{2}(y_{3}-y_{1})+x_{3}(y_{1}-y_{2}) \right ]"

"=\frac{1}{2}\left [ 1(-3+4)-4(-4-3)+5(3+3) \right ]"

"=\frac{1}{2}\left [ 1+28+30 \right ]
=\frac{59}{2}" square units